OpenStudy (anonymous):

Help please ! Find the rectangular coordinates of the point with the polar coordinates (-3, 5pi/3)

OpenStudy (anonymous):

Is it (-3/2, 3/2) ?

OpenStudy (anonymous):

\[x=r \cos \theta ,y=r \sin \theta \] \[r=-3,\theta =\frac{ 5 \pi }{ 3 }\] x co-ordinate is correct. calculate y co-ordinate again.

OpenStudy (superdavesuper):

the general eqn to convert from polar coordinates (r, theta) to rect coord (x,y) is: x=rcostheta and y=rsintheta

OpenStudy (anonymous):

is it (-3/2, 3pi3/2)?

OpenStudy (superdavesuper):

nope sorry: y=(-3)sin(5pi/3)

OpenStudy (anonymous):

That's not one of my options =\

OpenStudy (anonymous):

\[\sin \frac{ 5 \pi }{ 3 }=\sin \left( 2 \pi-\frac{ \pi }{ 3 } \right)=-\sin \frac{ \pi }{ 3 }=-\frac{ \sqrt{3} }{ 2 }\] now you can complete.

OpenStudy (anonymous):

Those are the options Im given

OpenStudy (superdavesuper):

y = \frac{ 3*\sqrt{3} }{ 2 }\] is not one of the option?